OpenStudy (anonymous):

3n+{2n-[-(3n+6)]+3n}

OpenStudy (anonymous):

11n+6

OpenStudy (anonymous):

{6 + 11 n}

OpenStudy (anonymous):

another answer pls

OpenStudy (anonymous):

3n+{2n-[-(3n+6)]+3n} = 3n+{2n +3n +6 +3n} = 3n+8n +6 = 11n + 6

OpenStudy (lalaly):

-3n-6 2n-(-3n-6)+3n 2n+3n+6+3n 8n-6 11n+6